Subshilava:

Keith is attempting to prove the Pythagorean Theorem by using similar triangles. After Keith has set up the appropriate ratios and solved, he ended up with two equations: a ^ 2 = cd b ^ 2 = ce What must Keith do next in order to complete the proof? A. Subtract the two equations. B. Multiply the two equations. C. Add the two equations. OD Divide the two equations.

Subshilava:

1 attachment
Subshilava:

@subshilava wrote:
Idk why every time I upload itโ€™s lags and uploads 25 images ๐Ÿ˜

toga:

nTo complete the proof of the Pythagorean Theorem using similar triangles, Keith needs to add the two equations a^2=cd and b^2=ce. This will give him a^2 + b^2 = cd + ce. From here, he can factor out the common term c and obtain a^2 + b^2 = c(d+e). Since d+e is equal to f (the length of the hypotenuse squared), he has shown that a^2 + b^2 = f, proving the Pythagorean Theorem.